The pKa of chloroacetic acid is 2.9. If an organic chemist has a 0.10 M aqueous solution of chloroacetic acid at pH 2.4, what percentage of molecules/ions will be in the basic form?
Answer #1

Let's denote the acid by HA and proceed to set up an ICE table with equilibrium dissociation of α. Also at pH = 2.4, there will be an initial H+ concentration of 3.98 * 10-3 M

                       HA →              H+         +      A-

initial conc:      0.1                            3.98 * 10-3                   0

equilibrium conc:      0.1(1-α)           3.98 * 10-3 + 0.1α              0.1α

We have pKa = 2.9

-log(Ka) = 2.9

Ka = 1.26 * 10-3

From the equation, Ka = [H+] * [A-] / [HA]

1.26 * 10-3 = (3.98 * 10-3 + 0.1α )* 0.1α / 0.1(1-α)     

Since α<<1, we assume 1-α = 1

Solving the equation, we have: α = 0.094

Since this is the fraction of acid that has dissociated, we can say that % of Base form = 100 * α= 9.4%